OpenStudy (anonymous):

A 60 kg physics student is working to pay his tuition fees by performing in a travelling carnival. He rides a motorcycle inside a hollow steel sphere of radius 15m. The surface of the sphere is full of holes so the audience can see in. After gaining sufficient speed, he travels in a vertical circle. What minimum velocity must he have at the top of the circle if the tires of the motorcycle are not to lose contact with the steel sphere?

OpenStudy (anonymous):

attempt to the solution: So: 9.8 = v^2 / 15 So: v = square root ( 15 * 9.8 ) v = 12.124 m/s is this correct?

OpenStudy (anonymous):

\[V=\sqrt{rg}\] Yes u r correct

OpenStudy (anonymous):

yey... thanks! x)

OpenStudy (anonymous):

Welcome..)